OpenStudy (anonymous):

Solve. x^2 = 6x – 6 x = 3 ± sqrt.3 x = 6 ± sqrt.3 x = 3 ± sqrt.30 x = 3 ± sqrt.6

OpenStudy (campbell_st):

rewrite the equation will all terms of the left. the general form of a quadratic is ax^2 + bx + c = 0 and the general quadratic formula is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] you need to identify a, b, c and then substitute and evaluate

OpenStudy (anonymous):

I substituted the terms into the quadratic formula but I have no idea what to do from there or how to go about solving it..

OpenStudy (campbell_st):

you just need to evaluate it...

OpenStudy (anonymous):

I got -6x+/- sqrt.6^2-4x^26/ 2x^2

OpenStudy (campbell_st):

well you need to know.... a,b and c are coefficients... do you know that term...?

OpenStudy (anonymous):

yea i herd it before..

OpenStudy (campbell_st):

do you know what it means...?

OpenStudy (campbell_st):

as an example if you have \[5x^2\] 5 is the coefficient and x is the pronumeral

OpenStudy (campbell_st):

so you need to identify the values a, b, c the coefficients

OpenStudy (anonymous):

ok so a would be 1, b would be 6 and c would be -6?

OpenStudy (campbell_st):

no... you need to rewrite the equation so it is equal to zero \[x^2 - 6x + 6 = 0\] then identify a, b, c..... substitute and evaluate

OpenStudy (anonymous):

so a=x^2, b=6x and c=6?

OpenStudy (anonymous):

the anwer is 3 +/- sqrt of 3

OpenStudy (campbell_st):

no a = 1, b = -6 and c = 6 your answer is correct.